### v1

# Evaluating time-specific treatment effects in matched-pairs studies

Abstract
This study develops methods for evaluating a treatment effect on a time-to-event outcome in matched-pair studies. While most methods for paired right-censored outcomes allow determining an overall treatment effect over the course of follow-up, they generally lack in providing detailed insights into how the effect changes over time. To address this gap, we propose time-specific and overall tests for paired right-censored outcomes under randomization inference. We further extend our tests to matched observational studies by developing corresponding sensitivity analysis methods to take into account departures from randomization. Simulations demonstrate the robustness of our approach against various non-proportional hazards alternatives, including a crossing survival curves scenario. We demonstrate the application of our methods using a matched observational study from the Korean Longitudinal Study of Aging (KLoSA) data, focusing on the effect of social engagement on survival.
1.   Introduction
1.1.   Motivating Example: The Effect of Social Engagement on Mortality
Kim et al. (2016) analyzed the Korean Longitudinal Study of Aging (KLoSA) cohort, which included individuals aged 45 years or older in 2006, to determine whether social engagement prolongs life in a late-middle-aged population. They found that active social engagement was significantly associated with a reduced risk of all-cause mortality using Cox proportional hazards models. Specifically, the hazard ratio for all-cause mortality was 1.84 times higher for those with the lowest level of social engagement compared to those with the highest level. Although the effect was significant when comparing survival functions between the high and low social engagement groups, the impact of active social engagement may change over the duration of a study. This variability over time in the effect of a treatment is known as time-varying treatment effects. For instance, it is reasonable to assume that the effect for an individual aged 45 will not be the same at age 50 as it is at age 70. We aim to address the following question: (1) Does the treatment effect of active social engagement actually vary over time, and (2) If it does vary, at which point is the effect strongest?
We analyze the KLoSA cohort data from 2006 to 2022 using matching to adjust for differences in baseline characteristics. Details on these covariates and the matching process are provided in Section 5. Figure 1 shows the survival functions for 1474 matched treated-control pairs after adjustment. Notably, the survival difference between the treated and control groups appears to be more pronounced between 6-8 follow-up years, diminishing towards the end of the study period. However, merely observing these differences is not a scientific method to evaluate the effect. In addition, ad hoc analysis does not control for statistical significance. In this paper, we propose a systematic approach to examine a time-varying effect in a matched observational study using a statistical hypothesis testing framework.
Figure 1: Kaplan-Meier survival curves for socially active individuals and their matched inactive controls for 1474 pairs.
We are also interested in the overall effect throughout the study period. One might wonder whether examining the time-varying effects could improve our understanding of the overall effect of the treatment. Studies on effect modification suggest that even when investigating the average treatment effect (ATE), acknowledging the presence of effect modification can provide a deeper understanding of how treatment effects manifest between different groups (Lee et al., 2018). More importantly, this understanding can lead to statistically more robust results (Lee et al., 2018; Lee and Hsu, 2023). For example, when conducting a sensitivity analysis, the results become less sensitive to unmeasured confounders. We believe that time-varying effects, much like effect modification, can offer a similar level of statistical robustness by illustrating how treatment effects change over time.
1.2.   Our contribution: Examining Multiple Time Points Simultaneously
In our motivating example, the outcome is time to death, which may be right-censored. Previous research has developed several nonparametric tests for analyzing such data, which are effective in identifying whether a treatment has an effect over the entire follow-up period (O’Brien and Fleming, 1987; Akritas, 1992; Dallas and Rao, 2000; Dabrowska, 1989; Murray, 2001). However, these tests do not provide detailed information about when the effect occurs or how long it lasts. In this work, we develop a time-specific test at a certain time point $\tau$ using randomization inference and evaluate the effect of treatment within the period up to $\tau$. The test uses time-specific scores at a time point $\tau$ by quantifying the impact of each individual on the survival function. A straightforward way to obtain the scores is to use pseudo-observations (Andersen et al., 2003). Recent advancement in pseudo-observations in survival analysis allows us to develop theoretical properties that will be discussed in Section 3.
In addition, we propose a powerful approach for jointly evaluating multiple time-specific tests at various time points. These tests, through assessing the effect multiple times, are correlated because they involve the same individuals over time. To maintain appropriate significance levels, we apply a method that considers the joint distribution of the test statistics rather than adjusting each test individually. This approach allows us to assess both the time-specific and overall treatment effects more effectively. Existing methods that simply aggregate survival differences at observed times may miss significant effects that appear only briefly or vary over time. Our approach, focusing on multiple time points, is specifically designed to capture these nuances with more accuracy.
Lastly, while most methods for analyzing right-censored outcomes are validated in controlled trials or well-adjusted observational studies, concerns about unmeasured confounders remain. Thus, we develop sensitivity analysis techniques to address potential biases from unmeasured confounders, ensuring that our conclusions about the causal effects of social engagement on mortality are robust. Sensitivity analysis has been well-established in matched observational studies (Rosenbaum, 1987, 2002), yet there is a noticeable lack of research on time-specific effects. We address this gap by effectively controlling the dependencies between the test statistics. Our proposed method will evaluate how sensitive our results are to different assumptions about the level of unmeasured confounding.
2.   Notation and Review
2.1.   Notation: Randomization Inference for Paired Right-Censored Outcomes in Experiments
There are $I$ independent pairs, $i=1,\dots,I$, of two units, $j=1,2$, one treated indicated by $Z_{ij}=1$ and the other control indicated by $Z_{ij}=0$, with $Z_{i1}+Z_{i2}=1$ for each $i$. The pairs are matched for measured covariates, $\bm{x}_{i1}=\bm{x}_{i2}$, but may possibly differ in terms of an unmeasured covariate, $u_{i1}\neq u_{i2}$. Following Neyman (1923) and Rubin (1974), each unit $ij$ exhibits survival (event) time $s_{1ij}$ if treated or survival time $s_{0ij}$ under control, so the realized survival time for $ij$ is $S_{ij}=Z_{ij}s_{1ij}+(1-Z_{ij})s_{0ij}$ and the effect of the treatment, $s_{1ij}-s_{0ij}$, on $ij$ cannot be observed for any unit. Similarly, let $c_{1ij}$ and $c_{0ij}$ denote the censoring (follow-up) times for $ij$ under treatment and control, respectively. We assume that each unit’s realized censoring time is deterministic, unrelated to the treatment $Z_{ij}$, that is, $c_{0ij}=c_{1ij}=c_{ij}$ for each $ij$. The realized variables $S_{ij}$ and $c_{ij}$ may not be observed in practice, but we can observe the minimum of the realized survival and censoring times, that is, $Y_{ij}=\min(S_{ij},c_{ij})$, and the event indicator for whether the event occurs before censoring, that is $\Delta_{ij}=I(S_{ij}\leq c_{ij})$, for each $ij$. Write $\bm{Y}=(Y_{11},\dots,Y_{I2})^{T}$, $\bm{\Delta}=(\Delta_{11},\dots,\Delta_{I2})^{T}$ for the $2I$ dimensional vectors, with a similar notation for $\bm{u}$. For simplicity of notation, the treated and control units in each pair $i$ are identified by the variable $V_{i}=Z_{i1}-Z_{i2}$, with $V_{i}=1$ if $i1$ is treated, and $V_{i}=-1$ if $i2$ is treated. Write $\bm{V}=(V_{1},...,V_{I})^{T}$ for the $I$ dimensional vector of treatment assignments for all $I$ pairs. Let $\mathcal{V}$ be the set containing the $2^{I}$ possible values $\bm{v}$ of $\bm{V}$, so that $\bm{v}\in\mathcal{V}$ if each $v_{i}$ is $1$ or $-1$. Moreover, write =$\mathcal{F}=\{(s_{1ij},s_{0ij},c_{ij},x_{ij},u_{ij}),i=1,\dots,I,j=1,2\}$. Then, the sets $\mathcal{F}$ and $\mathcal{V}$ are fixed through the course of randomization inference (Fisher, 1935).
In a paired experiment, random assignment of the treatment within pairs ensures that $\Pr(V_{i}=1|\mathcal{F},\mathcal{V})=\Pr(V_{i}=-1|\mathcal{F},\mathcal{V})=1/2$, $i=1,\dots,I$, so $\Pr(\bm{V}=\bm{v}|\mathcal{F},\mathcal{V})=1/2^{I}$ for each $\bm{v}\in\mathcal{V}$. For testing the sharp null hypothesis,
$H_{0}:s_{0ij}=s_{1ij}\ \text{for all}\ ij,$
given a test statistic $T(\bm{V},\bm{Y},\bm{\Delta})$ and its observed value $t$, the randomization $p$-value is therefore $\Pr(T\geq t|\mathcal{F},\mathcal{V})=|\{\bm{v}\in\mathcal{V}:T(\bm{v},\bm{Y},\bm{\delta})\geq t\}|/2^{I}$ because $Y_{ij}=\min(s_{0ij},c_{ij})$ and $\Delta_{ij}=I(s_{0ij}\leq c_{ij})$ are fixed under $H_{0}$ by conditioning on $\mathcal{F}$ and $\bm{V}$ is uniformly distributed on $\mathcal{V}$. Let $q_{ij}$ be a score assigned to unit $ij$ that depends only on $(\bm{Y},\bm{\Delta})$. For instance, $q_{ij}$ might be the log-rank score (Mantel, 1966) defined by $\hat{H}(Y_{ij})-\Delta_{ij}$, where $\hat{H}(a)=\sum_{k:Y_{k}\leq a}m_{k}/n_{k}$, $Y_{k}$ is an observed time among all $2I$ observations, $n_{k}$ is the number at risk at $Y_{k}$ and $m_{k}$ is the number of events that happened at $Y_{k}$. Alternatively, $q_{ij}$ might be the Prentice-Wilcoxon score (Kalbfleisch and Prentice, 1980) defined by $1-\tilde{S}(Y_{ij})-\Delta_{ij}\tilde{S}(Y_{ij})$, where $\tilde{S}(a)=\prod_{k:Y_{k}\leq a}(n_{k}-m_{k}+1)/(n_{k}+1)$. In O’Brien and Fleming (1987) and Dallas and Rao (2000), the resulting test statistics are equivalent in form to $T=\sum_{i=1}^{I}(q_{i1}-q_{i2})V_{i}=\sum_{i=1}^{I}d_{i}V_{i}$ with $d_{i}=q_{i1}-q_{i2}$. In this statistic, the $q_{ij}$’s are fixed under $H_{0}$ by conditioning on $\mathcal{F}$ and the $V_{i}$’s as the only stochastic quantities. In particular, pairs with tied scores $q_{i1}=q_{i2}$ have $d_{i}V_{i}=0$ regardless of the value of $V_{i}$, so they do not contribute any stochastic component to the statistic.
2.2.   Sensitivity Analysis in Observational Studies
In an observational study, failure to match on the unmeasured covariate $u_{ij}$ may lead to biased treatment assignments within pairs, $\Pr(\bm{V}=\bm{v}|\mathcal{F},\mathcal{V})\neq 1/2^{I}$. A simple model for sensitivity to departures from random treatment assignments assumes that the distribution of $Z_{ij}$ given measured covariates $\bm{x}_{ij}$ and a hypothetical unmeasured covariate $u_{ij}$ follows a logistic model,
$\log\frac{\Pr(Z_{ij}=1|\bm{x}_{ij},u_{ij})}{\Pr(Z_{ij}=0|\bm{x}_{ij},u_{ij})}=\kappa(\bm{x}_{ij})+\gamma u_{ij},$ (1)
where $\kappa(\cdot)$ is an arbitrary function, $\gamma\geq 0$ (or $\Gamma=\exp(\gamma)\geq 1$ equivalently) is a sensitivity parameter, and $u_{ij}\in[0,1]$ is normalized to make $\gamma$ more interpretable (Rosenbaum, 1987). The logistic model (1) precisely says that two units in pair $i$ will have odds of treatment that differ by the factor $1/\Gamma\leq\exp\{\gamma(u_{i1}-u_{i2})\}\leq\Gamma$, so the sensitivity parameter $\Gamma$ quantifies unmeasured confounding after matching on $\bm{x}_{ij}$.
Write $w_{i}=u_{i1}-u_{i2}$ and $\bm{w}=(w_{1},...w_{I})^{T}$. It is not difficult to verify that the logit model (1) is the same as assuming that, for each $\bm{v}\in\mathcal{V}$,
$\Pr(\bm{V}=\bm{v}|\mathcal{F},\mathcal{V})=\prod_{i=1}^{I}\frac{\exp(\frac{1}{2}\gamma v_{i}w_{i})}{\exp(\frac{1}{2}\gamma w_{i})+\exp(-\frac{1}{2}\gamma w_{i})}=\frac{\exp(\frac{1}{2}\gamma\bm{v}^{T}\bm{w})}{\sum_{\bm{b}\in\mathcal{V}}\exp(\frac{1}{2}\gamma\bm{b}^{T}\bm{w})},\bm{w}\in[-1,1]^{I}.$ (2)
If $\Gamma=1$, then the model (2) becomes the randomization distribution $\Pr(\bm{V}=\bm{v}|\mathcal{F},\mathcal{V})=1/2^{I}$, whereas if $\Gamma>1$, then the probability $\Pr(\bm{V}=\bm{v}|\mathcal{F},\mathcal{V})$ is unknown because $\bm{w}$ is unobserved, but the extent of the deviation from the randomization distribution is controlled by the value of $\Gamma$. Although the model (2) cannot be used for the direct adjustment of the $p$-value $\Pr(T\geq t|\mathcal{F},\mathcal{V})$ to take account of bias due to $\bm{u}$, the “worst-case $p$-value” defined as the largest $p$-value over all possible allocations of $\bm{w}\in[-1,1]^{I}$ can be found. Define $w_{i}^{+}=\text{sgn}(d_{i})$, where $\text{sgn}(a)=1$ if $a$ is positive, -1 if $a$ is negative, and 0 if $a$ is zero. Write $\bm{w}^{+}=(w_{1}^{+},\dots,w_{I}^{+})^{T}$. Let $T^{+}$ denote the distribution of $T=\sum_{i=1}^{I}d_{i}V_{i}$ when $\bm{w}$ equals $\bm{w}^{+}$. Then, under the model (2) and $H_{0}$, $T$ satisfies
$\Pr(T\geq t|\mathcal{F},\mathcal{V})\leq\Pr(T^{+}\geq t|\mathcal{F},\mathcal{V})\ \forall\ \bm{w}\in[-1,1]^{I},$
and as $I\rightarrow\infty$, the upper bound is approximated by
$\Pr(T^{+}\geq t|\mathcal{F},\mathcal{V})\overset{\cdot}{=}1-\Phi((t-\mu^{+})/\sigma^{+}),$
where $\mu^{+}$ and $\sigma^{+}$ are conditional mean and standard devition of $T^{+}$ given $(\mathcal{F},\mathcal{V})$ under the model (2) and $H_{0}$, and $\Phi(\cdot)$ is the standard normal cumulative distribution. The approximation holds when $\max_{i\in\{1,\dots,I\}}d_{i}^{2}/\sum_{i=1}^{I}d_{i}^{2}\to 0$ as $I\to\infty$ (Rosenbaum, 2007). For $\Gamma>1$, the upper bound is sharp, attained at $\bm{w}=\bm{w}^{+}$, and cannot be improved without further information about $\bm{u}$. This upper bound can be examined by considering several values of $\Gamma\geq 1$. The magnitude of $\Gamma$ indicates unmeasured bias that would need to be present to produce a worst-case $p$-value above a significance level $\alpha$. In practice, for a significance level of $\alpha$, we compute the (truncated) sensitivity value that is the smallest $\Gamma\geq 1$ such that the upper bound is not significant. This value serves as an indicator of the robustness of the analysis against unmeasured bias.
2.3.   The Power of a Sensitivity Analysis and Design Sensitivity
Suppose there is a treatment effect meaning that $H_{0}$ is false. Also, suppose matching for $\bm{x}_{ij}$ successfully eliminates overt bias in an observational study. If there is no unmeasured bias, $\Pr(\bm{V}=\bm{v}|\mathcal{F},\mathcal{V})=1/2^{I}$ for each $\bm{v}\in\mathcal{V}$ holds. If we were in situations where the treatment effect is non-null without unmeasured bias, called favorable situations, we could not be certain of this from the observed data. The best we could hope to say is that the results are insensitive to moderate biases $\Gamma$. The power of an $\alpha$ level sensitivity analysis is the probability that the worst-case $p$-value $\Pr(T^{+}\geq t|\mathcal{F},\mathcal{V})$ will be less than or equal to $\alpha$, computed in a given favorable situation (Rosenbaum, 2004). For a given model for generating $(\mathcal{F},\mathcal{V})$ with a treatment effect and without bias from $\bm{u}$, there typically exists a value $\tilde{\Gamma}$ called the design sensitivity such that as $I\rightarrow\infty$, the power of a sensitivity analysis tends to $1$ if the analysis is performed with $\Gamma<\tilde{\Gamma}$ and to 00 if performed with $\Gamma>\tilde{\Gamma}$. That is, in infinitely large sample sizes, the test can distinguish the model from biases smaller than $\tilde{\Gamma}$ but not from biases larger than $\tilde{\Gamma}$. See Rosenbaum (2004) and Rosenbaum (2010) for detailed discussions on design sensitivity, and see Zhao (2019) for insights into the relationship between sensitivity value and design sensitivity.
2.4.   Pseudo-observations
Though any time-specific scores can be used in our proposed method that will be introduced in Section 3, we introduce pseudo-observations (Andersen et al., 2003) in this subsection. Pseudo-observations have been adopted in a series of causal survival studies mainly for regression modeling of a survival summary, such as the survival function, the cause-specific cumulative incidence, or the restricted mean survival; see, for example, Kjaersgaard and Parner (2016) and Andersen et al. (2017). Let $\theta=E(f(S_{ij}))$ be a survival summary and $\hat{\theta}$ an estimate of $\theta$ based on all $2I=N$ observations. Given this estimate, the jackknife method generates pseudo-observations by leaving out one observation at a time. Specifically, the pseudo-observation of $\hat{\theta}$ for unit $ij$ is defined as
$\hat{\theta}_{ij}=N\hat{\theta}-(N-1)\hat{\theta}^{-ij},$
where $\hat{\theta}^{-ij}$ is the estimate obtained by leaving unit $ij$ out of the sample and recalculating $\hat{\theta}$ based on $N-1$ observations $i^{\prime}j^{\prime}\neq ij$. Intuitively, the pseudo-observation $\hat{\theta}_{ij}$ can be thought of as the contribution of unit $ij$ to the overall estimate $\hat{\theta}$ based on all $N$ observations.
Example 1 (Survival function)
Suppose the survival summary $\theta$ of interest is the survival function $S(\tau)$ at time $\tau>0$. In this example, the function $f(\cdot)$ is given by $f(S_{ij})$ = $I(S_{ij}>\tau)$. Using the Kaplan-Meier estimate $\hat{S}(\tau)=\prod_{k:Y_{k}\leq\tau}(1-m_{k}/n_{k})$, the pseudo-observation for unit $ij$ is defined as
$\hat{S}_{ij}(\tau)=N\hat{S}(\tau)-(N-1)\hat{S}^{-ij}(\tau),$
where $\hat{S}^{-ij}(\tau)$ is the Kaplan-Meier estimate of $S(\tau)$ based on $N-1$ observations $i^{\prime}j^{\prime}\neq ij$.
Example 2 (Cause-specific cumulative incidence)
Suppose the survival summary $\theta$ of interest is the cumulative incidence $F_{r}(\tau)$ due to cause $r$ at time $\tau>0$. In this case, the function $f(\cdot)$ is given by $f(S_{ij})$ = $I(S_{ij}\leq\tau,D_{ij}=r)$, where $D_{ij}\in\{1,\dots,R\}$ indicates which competing risk caused the event. The estimate of $F_{r}(\tau)$ may be the Aalen-Johansen estimate (Aalen and Johansen, 1978) $\hat{F}_{r}(\tau)=\sum_{l:Y_{l}\leq\tau}(m_{l}^{(r)}/n_{l})\prod_{k:Y_{k}<Y_{l}}(1-m_{k}/n_{k})$, where $m_{l}^{(r)}$ is the number of events due to cause $r$ at $Y_{l}$. Using $\hat{F}_{r}(\tau)$, the pseudo-observation for unit $ij$ is defined as
$\hat{F}_{r,ij}(\tau)=N\hat{F}_{r}(\tau)-(N-1)\hat{F}_{r}^{-ij}(\tau),$
where $\hat{F}_{-ij}(\tau)$ is the Aalen-Johansen estimate of $F_{r}(\tau)$ based on $N-1$ observations $i^{\prime}j^{\prime}\neq ij$.
The pseudo-observations defined in Example 1 and Example 2 can be calculated for any $\tau>0$, but their values only change at event times, being constant in between. In subsection 3.3, we will discuss a desirable large-sample property of these pseudo-observations. This property will be needed to derive some auxiliary theorems.
3.   Tests for the Time-Specific Treatment Effect
3.1.   Time-Specific Null Hypothesis
In this subsection, we aim to assess a treatment effect within a specific duration $(0,\tau]$. Specifically, we are interested in testing the partially sharp null hypothesis,
$H_{0}(\tau):\min(s_{0ij},\tau)=\min(s_{1ij},\tau)\ \text{for any}\ ij,$
which focuses only on the treatment’s effect up to the time $\tau$, ignoring any difference in survival beyond that time. We define $q_{ij}(\tau)$ to be a score for unit $ij$. Although this score can be chosen in many different ways, for the remainder of this paper, we will use the pseudo-observation based on the Kaplan-Meier estimate $\hat{S}(\tau)$ for unit $ij$, denoted as $q_{ij}(\tau)=\hat{S}_{ij}(\tau)$. Note that $q_{ij}(\tau),i=1,\dots,I,j=1,2$ are fixed under $H_{0}(\tau)$ by conditioning on $\mathcal{F}$ despite that $H_{0}(\tau)$ is not fully sharp. To see a concrete example, let $\tau>1$ and consider $I=1$ pair of two uncensored units. If unit $11$ exhibits survival time $s_{111}=\tau+5$ if treated or survival time $s_{011}=\tau+1$ under control and unit $12$ exhibits $s_{112}=s_{012}=\tau-1$ whether treated or control, then $H_{0}(\tau)$ holds for this pair because $\min(s_{011},\tau)=\min(s_{111},\tau)=\tau$ and $\min(s_{012},\tau)=\min(s_{112},\tau)=\tau-1$ but $H_{0}$ does not due to $s_{111}\neq s_{011}$. Yet, $q_{11}(\tau)$ and $q_{12}(\tau)$ are fixed at $1$ and 00 regardless of the value of $V_{1}$ by conditioning on $\mathcal{F}$ since neither the overall estimate $\hat{S}(\tau)$ nor the jackknife estimates $\hat{S}_{-11}(\tau)$ and $\hat{S}_{-12}(\tau)$ are affected by the the discrepancy between $s_{111}=\tau+5$ and $s_{011}=\tau+1$.
In parallel to the notation in Subsection 2.1, define $T_{\tau}=\sum_{i=1}^{I}(q_{i1}(\tau)-q_{i2}(\tau))V_{i}=\sum_{i=1}^{I}d_{i}(\tau)V_{i}$ with $d_{i}(\tau)=q_{i1}(\tau)-q_{i2}(\tau)$. Let $t_{\tau}$ be the observed value of $T_{\tau}$. Write $w_{\tau,i}^{+}=\text{sgn}(d_{i}(\tau))$ and $\bm{w}^{+}_{\tau}=(w_{\tau,1}^{+},\dots,w_{\tau,I}^{+})^{T}$. Moreover, let $T^{+}_{\tau}$ denote the distribution of $T_{\tau}$ when $\bm{w}=\bm{w}^{+}_{\tau}$. By the arguments in subsection 2.2, under the model (2) and $H_{0}(\tau)$, $T^{+}_{\tau}$ stochastically dominates $T_{\tau}$ in such a way that
$\Pr(T_{\tau}\geq t_{\tau}|\mathcal{F},\mathcal{V})\leq\Pr(T^{+}_{\tau}\geq t_{\tau}|\mathcal{F},\mathcal{V})\ \text{for all}\ \bm{w}\in[-,1,1]^{I},$ (3)
and as $I\rightarrow\infty$ the upper bound in (3) is approximated by $\Pr(T^{+}_{\tau}\geq t_{\tau}|\mathcal{F},\mathcal{V})\overset{\cdot}{=}1-\Phi\left((t-\mu_{\tau}^{+})/\sigma_{\tau}^{+}\right),$ where $\mu_{\tau}^{+}$ and $\sigma_{\tau}^{+}$ denote the conditional mean and standard deviation of $T^{+}_{\tau}$ given $(\mathcal{F},\mathcal{V})$ under the model (2) and $H_{0}(\tau)$. Specifically, $\mu_{\tau}^{+}=[(\Gamma-1)/(1+\Gamma)]\sum_{i=1}^{I}|d_{i}(\tau)|$ and $(\sigma_{\tau}^{+})^{2}=[4\Gamma/(1+\Gamma)^{2}]\sum_{i=1}^{I}d_{i}^{2}(\tau)$. We note that an exact randomization test with $\Gamma=1$ or an exact sensitivity analysis with $\Gamma>1$ could also be obtained by considering all $2^{I}$ possible values of $T^{+}_{\tau}$.
3.2.   Test of No Treatment Effect At All
We now address the problem of testing the sharp null hypothesis $H_{0}$ of absolutely no effect. Although we can test $H_{0}$ directly using the scores discussed in Section 2.1, we choose to indirectly test it through the intersection of $L$ overlapping hypotheses $H_{0}(\tau)$,
$H_{0}^{*}=\cap_{l=1}^{L}H_{0}(\tau_{l}).$
The null hypothesis $H_{0}^{*}$ is not identical to $H_{0}$, but includes it. As such, rejecting $H_{0}^{*}$ is more challenging if there is an effect. When $\tau$ is sufficiently large enough so that all survival times are observed, or simply $\tau=\infty$, $H_{0}(\tau)$ should be the same as $H_{0}$. By definition, rejecting at least one of $H_{0}(\tau_{l})$ leads to the rejection of $H_{0}^{*}$, and consequently $H_{0}$. It might appear counter-intuitive to test $H_{0}^{*}$ rather than $H_{0}$. Existing methods that test $H_{0}$ typically aggregate differences between treated and control survival functions at each observed time point. However, if these differences vary with time points, then any significant difference at a single time point may be diluted when aggregated. A test that directly uses $H_{0}$ may fail to detect this time-specific difference. In contrast, our proposed hypothesis, which considers multiple time points, is more likely to capture any time-specific effects. We will demonstrate that the proposed test can be more powerful when the magnitude of a treatment effect varies across different time points.
The overall null hypothesis $H_{0}^{*}$ is tested by performing $L\geq 2$ time-specific tests using $T_{\tau}$ and adjusting for multiple testing error exploiting their joint distribution of $T_{\tau_{l}}$. We propose a novel approach to find this joint distribution and extend it to finding the joint distribution of $T_{\tau_{l}}^{+}$, which enables us to find the $p$-value upper bound even when there is unmeasured bias or $\Gamma>1$. Though many studies on combining dependent tests have focused on developing methods for combining $p$-values, our approach aims to combine the statistics by computing the dependency structure directly. This approach has been successful in testing the overall null hypothesis, as well as the specific null hypothesis (Rosenbaum, 2012; Lee et al., 2018).
Define $\mathcal{T}$ to be the set of $L$ different time points $\tau_{1}<\ldots<\tau_{L}$ such that the numbers of events expected between them are sufficiently large. Assume $H_{0}$ is true so that all related scores $q_{ij}(\tau_{l}),i=1,\dots,I,j=1,2,\tau\in\mathcal{T}$ are fixed by conditioning on $\mathcal{F}$. In the absence of bias from $\bm{u}$, it is easy to check that the conditional covariance of $T_{\tau_{k}}$ and $T_{\tau_{l}}$ given $(\mathcal{F},\mathcal{V})$ is $\sum_{i=1}^{I}d_{i}(\tau_{k})d_{i}(\tau_{l})=\sigma_{\tau_{k}\tau_{l}}$ for $\tau_{k},\tau_{l}\in\mathcal{T}$. Write $\sigma_{\tau_{l}}^{2}=\sigma_{\tau_{l}\tau_{l}}$ for simplicity. Let $\rho_{\tau_{k}\tau_{l}}=\sigma_{\tau_{k}\tau_{l}}/(\sigma_{\tau_{k}}\sigma_{\tau_{l}})$ and $\bm{\rho}$ be the $L\times L$ correlation matrix containing $\rho_{\tau_{k}\tau_{l}}$. To test the null hypothesis $H_{0}$, we propose using the maximum of all standardized statistics at all times points, $M_{\mathcal{T}}=\max_{\tau_{l}\in\mathcal{T}}T_{\tau_{l}}/\sigma_{\tau_{l}}$. Theorem 1 states a multidimensional central limit theorem for the joint distribution of $T_{\tau_{l}}/\sigma_{\tau_{l}}$, providing a way of conducting an approximate randomization test of $H_{0}$ using $M_{\mathcal{T}}$ as the test statistic. The testing procedure is as follows:
$\text{reject $H_{0}$ \, if }M_{\mathcal{T}}=\max_{\tau_{l}\in\mathcal{T}}T_{\tau_{l}}/\sigma_{\tau_{l}}\geq\kappa_{\alpha}\quad\text{ where }\Pr(M_{\mathcal{T}}<\kappa_{\alpha}|\mathcal{F},\mathcal{V})=1-\alpha.$ (4)
The critical value $\kappa_{\alpha}$ is obtained using the qmvnorm function in the mvtnorm package in R.
Theorem 1
Assume treatments were assigned at random within pairs to satisfy $\Gamma=1$. Under $H_{0}$, if
$\max_{\tau_{l}\in\mathcal{T}}\left\{\max_{i\in\{1,\dots,I\}}\frac{d_{i}^{2}(\tau_{l})}{\sum_{i=1}^{I}d_{i}^{2}(\tau_{l})}\right\}\rightarrow 0\quad\text{ as }\>I\rightarrow\infty,$
then the random vector $(T_{\tau_{1}}/\sigma_{\tau_{l}},\cdots,T_{\tau_{L}}/\sigma_{\tau_{L}})^{T}$ converges in distribution to the $L$-dimensional normal distribution $N_{L}(\bm{0},\bm{\rho})$ with expectation $\bm{0}$ and covariance matrix $\bm{\rho}$; in turn,
$\Pr(M_{\mathcal{T}}\geq m_{\mathcal{T}}|\mathcal{F},\mathcal{V})\overset{\cdot}{=}1-\Pr(\chi_{l}\leq m_{\mathcal{T}},\cdots,\chi_{L}\leq m_{\mathcal{T}})\quad\text{ as }\>I\to\infty,$ (5)
where $(\chi_{1},\dots,\chi_{L})^{T}$ is distributed as $N_{L}(\bm{0},\bm{\rho})$.
Theorem 1 can be further extended to a sensitivity analysis for a specific value of $\Gamma$. Recall that $T^{+}_{\tau_{l}}$ denotes the distribution of $T_{\tau_{l}}$ when $\bm{w}=\bm{w}^{+}_{\tau_{l}}$. Under the model and $H_{0}$, it follows from (3) that
$\begin{split}\textstyle\Pr(M_{\mathcal{T}}\geq m_{\mathcal{T}}|\mathcal{F},\mathcal{V})\leq\Pr(\max_{\tau_{l}\in\mathcal{T}}T^{+}_{\tau_{l}}/\sigma_{\tau_{l}}\geq m_{\mathcal{T}}|\mathcal{F},\mathcal{V})\ \text{for all}\ \bm{w}\in[-1,1]^{I}.\end{split}$ (6)
In the case of no censoring, the signs of the nonzero values among the $L$ differences $d_{i}(\tau_{1}),\dots,d_{i}(\tau_{L})$ are concordant for each $i$ due to the cumulative nature of the Kaplan-Meier estimator. This results in $\bm{w}_{\tau_{1}}^{+}=\dots=\bm{w}_{\tau_{L}}^{+}$ (forcing the value of $\bm{w}_{\tau_{1}}^{+}$ for any tied pair $i$ to meet the equalities). Consequently, $\max_{\bm{w}\in[-1,1]^{I}}\Pr(M_{\mathcal{T}}\geq m_{\mathcal{T}}|\mathcal{F},\mathcal{V})$ is attained when $\bm{w}=\bm{w}_{\tau_{1}}^{+}=\dots=\bm{w}_{\tau_{L}}^{+}$, meaning that inequality (6) is sharp. In contrast, when observations are censored, it may not be the case that $\bm{w}_{\tau_{1}}^{+}=\dots=\bm{w}_{\tau_{L}}^{+}$. For example, suppose we have five matched pairs with $\bm{Y}=(8.3,1.8,4.8,9.8,4.5,11.4,5.8,9.4,5.9,1.3)^{T}$ and $\bm{\Delta}=(1,1,1,1,1,0,0,1,1,1)^{T}$. Then, the specific treatment assignment $\bm{V}=(1,1,1,1,1)^{T}$ yields $d_{5}(1.3)=-1$ and $d_{5}(5.9)=1/5$ as the censored unit $41$ with $Y_{41}=5.8$ is excluded from the “at risk” set before $\tau=5.9$.
With censoring, the upper bound in (6) may be somewhat conservative but readily solved by a normal approximation as stated in Theorem 2. We first introduce the following notation to formulate Theorem 2. Under the model (2) and $H_{0}$, it is easily derived that the conditional covariance of $T_{\tau_{k}}^{+}$ and $T_{\tau_{l}}^{+}$ given $(\mathcal{F},\mathcal{V})$ is $\{4\Gamma/(1+\Gamma)^{2}\}\sum_{i=1}^{I}|d_{i}(\tau_{k})d_{i}(\tau_{l})|=\sigma_{\tau_{k}\tau_{l}}^{+}$ for $\tau_{k},\tau_{l}\in\mathcal{T}$. Notice that $\sigma_{\tau_{l}\tau_{l}}^{+}=\sigma_{\tau_{l}}^{+2}$. Write $\rho_{\tau_{k}\tau_{l}}^{+}=\sigma_{\tau_{k}\tau_{l}}^{+}/(\sigma_{\tau_{k}}^{+}\sigma_{\tau_{l}}^{+})$ and $\bm{\rho}^{+}$ for the $L\times L$ correlation matrix containing $\rho_{\tau_{k}\tau_{l}}^{+}$.
Theorem 2
Under the model (2) and $H_{0}$, if the condition in Theorem 1 holds, then as $I\to\infty$ the random vector $((T_{\tau_{1}}^{+}-\mu_{\tau_{1}}^{+})/\sigma_{\tau_{1}}^{+},\dots,(T_{\tau_{L}}^{+}-\mu_{\tau_{L}}^{+})/\sigma_{\tau_{L}}^{+})^{T}$ converges in distribution to the $L$-dimensional normal distribution $N_{L}(\bm{0},\bm{\rho}^{+})$ with expectation $\bm{0}$ and covariance matrix $\bm{\rho}^{+}$; in turn,
$\Pr(\max_{\tau_{l}\in\mathcal{T}}T^{+}_{\tau_{l}}/\sigma_{\tau_{l}}\geq m_{\mathcal{T}}|\mathcal{F},\mathcal{V})$
$\overset{\cdot}{=}1-\Pr\left(\chi_{1}<\frac{m_{\mathcal{T}}\sigma_{\tau_{1}}-\mu_{\tau_{1}}^{+}}{\sigma_{\tau_{1}}^{+}},\dots,\chi_{L}<\frac{m_{\mathcal{T}}\sigma_{\tau_{L}}-\mu_{\tau_{L}}^{+}}{\sigma_{\tau_{L}}^{+}}\right)\quad\text{as}\quad I\to\infty,$ (7)
where $(\chi_{1},\dots,\chi_{L})^{T}$ is distributed as the $L$-dimensional normal distribution $N_{L}(\bm{0},\bm{\rho}^{+})$ with expectation $\bm{0}$ and covariance matrix $\bm{\rho}^{+}$.
Note that the normal approximations (5) and (2) are adequate when the distributions $N_{L}(\bm{0},\bm{\rho})$ and $N_{L}(\bm{0},\bm{\rho}^{+})$ are non-degenerate. This condition will possibly be ensured if the $L$ analysis times in the set $\mathcal{T}$ are spaced enough, inhibiting strong correlations between the $L$ associated test statistics $T_{\tau_{l}}$. As we noted previously, an exact randomization test of $H_{0}$ at $\Gamma=1$ or a corresponding exact sensitivity analysis at $\Gamma>1$ could also be obtained using the permutational distribution of $M_{\mathcal{T}}$ in small samples.
3.3.   Formulas for Design Sensitivities
In this section, we derive design sensitivity formulas for the time-specific test in Section 3.1 and the overall test in Section 3.2 to evaluate their asymptotic performance in sensitivity analyses. Deriving the formulas requires an investigation into the large-sample properties of pseudo-observations. Lemma 1 says that, without any competing cause-of-event ($R=1$), as $N\to\infty$, the pseudo-observations of the Aalen-Johansen estimate $\hat{F}(\tau)=\hat{F}_{r}(\tau)$ of $F(\tau)$ are approximated by independent and identically distributed (i.i.d.) variables under random (or non-informative) censoring.
Lemma 1
Assume there is no competing cause-of-event and $S_{ij}\perp\!\!\!\perp c_{ij}$. Then, the pseudo-observations of the Aalen-Johansen estimate $\hat{F}(\tau)$ are represented as
$\begin{split}\hat{F}_{ij}(\tau)&=N\hat{F}(\tau)-(N-1)\hat{F}^{-ij}(\tau)\\ &=F(\tau)+\dot{\psi}(Y_{ij},\Delta_{ij};\tau)+o_{P}(1),i=1,\dots,I,j=1,2\end{split}$
where $\dot{\psi}(\cdot;\tau)$ is the first-order influence curve of the Aalen-Johansen functional.
Lemma 1 is derived using the second-order von Mises expansion of the Aalen-Johansen functional, as shown by Graw et al. (2009). As a consequence of Lemma 1, under the conditions in Lemma 1, as $N\to\infty$, the property of i.i.d. also holds for $q_{ij}(\tau)$, $i=1,\dots,I,j=1,2$ because the Aalen-Johansen estimate $\hat{F}(\tau)$ reduces to $1-\hat{S}(\tau)$ in the case of no competing cause-of-event (Andersen and Pohar Perme, 2010).
Consider whether $T_{\tau}$ would lead to rejection of $H_{0}(\tau)$ in a sensitivity analysis under the model (2). Using (5), as $I\rightarrow\infty$, the approximate worst-case $p$-value is less than or equal to $\alpha$ if
$\Phi^{-1}(1-\alpha)\leq\frac{\frac{\sum_{i=1}^{I}d_{i}(\tau)V_{i}/I}{\sqrt{\sum_{i=1}^{I}d_{i}^{2}(\tau)/I}}-\frac{\Gamma-1}{1+\Gamma}\frac{\sum_{i=1}^{I}|d_{i}(\tau)|/I}{\sqrt{\sum_{i=1}^{I}d_{i}^{2}(\tau)/I}}}{\sqrt{\frac{4\Gamma}{(1+\Gamma)^{2}}/I}}=D_{\tau,I}$
is satisfied. This means that $H_{0}(\tau)$ is rejected for all possible allocations of $\bm{w}\in[-1,1]^{I}$. The power of this $\alpha$ level sensitivity analysis asymptotically equals the probability $\Pr(D_{\tau,I}\geq\Phi^{-1}(1-\alpha))$ under a given favorable situation with a treatment effect and no bias from $\bm{u}$.
Theorem 3
Consider a particular data-generating model for $(\mathcal{F},\mathcal{V})$ with a treatment effect and without bias from $\bm{u}$. If the model satisfies the assumptions in Lemma 1 and if $\dot{\psi}(\cdot;\tau)$ is bounded, then for all $\alpha\in(0,1)$, as $I\rightarrow\infty$, $\Pr(D_{\tau,I}\geq\Phi^{-1}(1-\alpha))\rightarrow 1$ if $\Gamma<\tilde{\Gamma}_{\tau}$ and $\Pr(D_{\tau,I}\geq\Phi^{-1}(1-\alpha))\rightarrow 0$ if $\Gamma>\tilde{\Gamma}_{\tau}$, where the design sensitivity $\tilde{\Gamma}_{\tau}$ is
$\tilde{\Gamma}_{\tau}=\frac{E(|d_{i}(\tau)|)+E(d_{i}(\tau)V_{i})}{E(|d_{i}(\tau)|)-E(d_{i}(\tau)V_{i})}.$ (8)
We now examine the limiting sensitivity of the overall test using $M_{\mathcal{T}}$ to bias from $\bm{u}$ under the model (2). Using (2), as $I\rightarrow\infty$, $H_{0}$ would be rejected for all possible allocations of $\bm{w}\in[-,1,1]^{I}$ if
$1-\alpha\leq\Pr\left(\chi_{l}\leq\frac{{\displaystyle\max_{\tau_{l}\in\mathcal{T}}}\frac{\sum_{i=1}^{I}d_{i}(\tau_{l})V_{i}/I}{\sqrt{\sum_{i=1}^{I}d_{i}^{2}(\tau_{l})/I}}-\frac{\Gamma-1}{1+\Gamma}\frac{\sum_{i=1}^{I}|d_{i}(\tau_{l})|/I}{\sqrt{\sum_{i=1}^{I}d_{i}^{2}(\tau_{l})/I}}}{\sqrt{\frac{4\Gamma}{(1+\Gamma)^{2}}/I}},l=1,\dots,L\right)=\Lambda_{\mathcal{T},I}$
The probability $\Pr(\Lambda_{\mathcal{T},I}\geq 1-\alpha)$ under a given favorable situation for $\Gamma=1$ in this case is not necessarily the same as the power of a randomization test since the relation (6) is not always equality at $\Gamma=1$.
Theorem 4
Consider a particular data-generating model for $(\mathcal{F},\mathcal{V})$ with a treatment effect and without bias from $\bm{u}$. If the model satisfies the assumptions in Lemma 1 and if $\dot{\psi}(\cdot;\tau_{l})$ is bounded for each $\tau_{l}\in\mathcal{T}$, then for all $\alpha\in(0,1)$, as $I\to\infty$, the power $\Pr(\Lambda_{\mathcal{T},I}\geq 1-\alpha)\to 1$ if $\Gamma<\tilde{\Gamma}_{\mathcal{T}}$ and $\Pr(\Lambda_{\mathcal{T},I}\geq 1-\alpha)\to 0$ if $\Gamma>\tilde{\Gamma}_{\mathcal{T}}$, where the design sensitivity $\tilde{\Gamma}_{\mathcal{T}}$ is
$\tilde{\Gamma}_{\mathcal{T}}=\frac{{\displaystyle\max_{\tau_{l}\in\mathcal{T}}}\frac{E(|d_{i}(\tau_{l})|)}{\sqrt{E(d_{i}^{2}(\tau_{l}))}}+{\displaystyle\max_{\tau_{l}\in\mathcal{T}}}\frac{E(d_{i}(\tau_{l})V_{i})}{\sqrt{E(d_{i}^{2}(\tau_{l}))}}}{{\displaystyle\max_{\tau_{l}\in\mathcal{T}}}\frac{E(|d_{i}(\tau_{l})|)}{\sqrt{E(d_{i}^{2}(\tau_{l}))}}-{\displaystyle\max_{\tau_{l}\in\mathcal{T}}}\frac{E(d_{i}(\tau_{l})V_{i})}{\sqrt{E(d_{i}^{2}(\tau_{l}))}}}.$ (9)
3.4.   Test for Time-Specific Treatment Effects
When the overall null hypothesis $H_{0}$ is rejected, researchers are interested in determining when the effect occurred. In this section, we test the time-specific hypothesis $H_{0}(\tau_{l})$ using the closed testing procedure (Marcus et al., 1976). The testing procedure is simple yet effectively controls the family-wise type 1 error rate. We can reject $H_{0}(\tau_{l})$ at a significance level $\alpha$ if all possible intersection hypotheses involving $H_{0}(\tau_{l})$ are also rejected at this level. Consider the hypothesis $H_{0}^{A}=\cap_{\tau_{l}\in A}H_{0}(\tau_{l})$, where $A$ is a subset of $\mathcal{T}=\{\tau_{1},\ldots,\tau_{L}\}$. Denote $\mathcal{I}_{A}\subset\mathcal{T}$ as any subset containing $A$ and $\mathcal{C}_{A}$ as the collection of all possible subsets $\mathcal{I}_{A}$. For example, suppose $\mathcal{T}=\{\tau_{1},\tau_{2},\tau_{3}\}$ and the hypothesis $H_{0}(\tau_{3})$, that is, $A=\{\tau_{3}\}$. Thus, $\mathcal{I}_{\{\tau_{3}\}}$ can be any element of the collection $\mathcal{C}_{\{\tau_{3}\}}=\left\{\{\tau_{3}\},\{\tau_{1},\tau_{3}\},\{\tau_{2},\tau_{3}\},\{\tau_{1},\tau_{2},\tau_{3}\}\right\}$. For each $\mathcal{I}_{A}$, we test the intersection hypothesis $H_{0}^{\mathcal{I}_{A}}:=\cap_{\tau_{l}\in\mathcal{I}_{A}}H_{0}(\tau_{l})$ and denote its $p$-value by $p(\mathcal{I}_{A})$. If the maximum $p$-value across all $\mathcal{I}_{A}$ is less than or equal to $\alpha$ (i.e., $\max_{\mathcal{I}_{A}}p(\mathcal{I}_{A})\leq\alpha$), then $H_{0}^{A}$ is rejected. This method can also be adapted for settings where $\Gamma>1$.
Computationally, to test $H_{0}^{\mathcal{I}_{A}}$, we derive a new test statistic $M_{\mathcal{I}_{A}}=\max_{\tau_{l}\in\mathcal{I}_{A}}T_{\tau_{l}}/\sigma_{\tau_{l}}$. The $p$-value $p(\mathcal{I}_{A})$ can be computed by using the sub-matrix $\rho_{\mathcal{I}_{A}}$ that selects columns associated with $\mathcal{I}_{A}$. Since $\mathcal{I}_{A}\subset\mathcal{T}$, the multiple testing burden for testing $H_{0}^{\mathcal{I}_{A}}$ is less severe than for $H_{0}$. Although not guaranteed, in the case where $M_{A}=M_{\mathcal{T}}$, $p(\mathcal{I}_{A})\leq\alpha$ when $p(\mathcal{T})\leq\alpha$. We will demonstrate in Section 5 how this closed testing procedure can be implemented.
4.   Simulations
4.1.   Size and Power Computation
We perform Monte Carlo simulations to investigate the sizes and powers of the time-specific tests of $H_{0}(\tau)$ using $T_{\tau}$ and the overall test of $H_{0}$ using $M_{\mathcal{T}}$. For this simulation study, we assume no unmeasured confounding ($\Gamma=1$). To evaluate the performance of our proposed methods, we also compare $M_{\mathcal{T}}$ with the Prentice-Wilcoxon test statistic (O’Brien and Fleming, 1987), which has been widely applied to matched pair data (Rosenbaum, 2002, 2010).
To generate survival and censoring times, we consider the hazard models $h_{s}(\tau|x_{i},z)=\lambda\exp(x_{i}+\eta(\tau,z))$ for $s_{zij}$, $z=0,1$ and $h_{c}(\tau|x_{i})=(\lambda/b)\exp(x_{i})$ for $c_{ij}$, where $x_{i}\sim N(0,1)$ is the common variate for each pair, $\eta(\cdot)$ is an arbitrary function, $\lambda=0.2$, and $b>1$ is a constant for adjusting the censoring rate. We set the administrative censoring time (i.e., the maximum censoring time) at 5 to ensure that no observed time exceeds this limit. We generate $2000$ replications consisting of $I=500$ independent paired sets of censored survival times under each of the following five scenarios: 1) $\eta(\tau,z)=0$, 2) $\eta(\tau,z)=-0.4z$, 3) $\eta(\tau,z)=(0.1\tau-0.5)z$, 4) $\eta(\tau,z)=(0.3\tau-0.6)z$, and 5) $\eta(\tau,z)=(0.15-0.14z)\tau$. The specifics of the data-generating processes follow (Austin, 2012). The first scenario indicates no effect of $z$ on survival (referred to as No effect). The second scenario corresponds to a proportional hazards model (referred to as PH). The third scenario (referred to as Early-div) and the fourth scenario (referred to as Crossing) both lead the survival curves for $z=0,1$ to diverge early on in favor of $z=1$, but the latter allows the curves to cross before the administrative censoring time $5$. On the contrary, the fifth scenario leads the survival curves for $z=0,1$ to exhibit late divergence in favor of $z=1$ (referred to as Late-div). For each of these five scenarios, $b$ is chosen separately to achieve moderate (approximately $25\%$) non-administrative censoring. In each resulting pair, one is randomly labeled $Z_{ij}=1$ and the other is labeled $Z_{ij}=0$, with the realized variable $S_{ij}$ set equal $s_{1ij}$ or $s_{0ij}$ accordingly. Figure 2 shows the Kaplan-Meier survival curves for the treated ($Z_{ij}=1$) and control ($Z_{ij}=0$) groups for a single sample of $I=500$ matched pairs simulated under the respective scenario.
Figure 2: Survival configurations used in our simulations. Each plot shows the Kaplan-Meier survival curves for the groups of $Z_{ij}=1$ and $Z_{ij}=0$ for a single sample of $I=500$ pairs simulated under respective scenario.
For illustration, we test $H_{0}(\tau)$ at $\tau=1,2,3,4,\ \text{and}\ 5$ using $T_{\tau}$, and test $H_{0}$ using $M_{\mathcal{T}}$ with $\mathcal{T}=\{1,2,3,4,5\}$. Additionally, we consider the Prentice-Wilcoxon test statistic defined as $PW=\sum_{i=1}^{I}d_{i}V_{i}$, where $d_{i}=q_{i1}-q_{i2}$ and $q_{ij}$ is the Prentice-Wilcoxon score defined in Section 2.1. We use the pairwise univariate censoring assumed in O’Brien and Fleming (1987) to compute the $q_{ij}$’s. For ease of computation, we apply normal approximation to obtain randomization $p$-values.
Table 1 summarizes size and power of tests. Under the No effect scenario, all tests maintain their empirical sizes close to $\alpha=0.05$. Under the PH scenario, $T_{\tau+1}$ exhibits higher power relative to $T_{1}\dots,T_{\tau}$ despite that the scenario assumes a constant effect over time. This is not surprising since the number of informative pairs with $d_{i}(\tau)\neq 0$ increases as time elapses, which also explains why $T_{1}$ has slightly lower power than $T_{2}$ under the Early-div scenario. Other than this, the empirical powers of the time-specific tests concordantly reflect the distinct time-varying effects assumed in Early-div, Crossing, and Late-div scenarios. The general pattern of the results of $PW$ agrees with those presented in Woolson and O’Gorman (1992); Dallas and Rao (2000); Sun and Sherman (1996); Lee and Wang (2003), suffering from power loss in the Late-div and Crossing scenarios. Overall, $M_{\mathcal{T}}$ outperforms $PW$, demonstrating robustness across different non-proportional hazards scenarios. In particular, the power using $M_{\mathcal{T}}$ is considerably higher compared to $PW$ in the Crossing and Late-div scenarios.
Table 1: Proportions of randomization $p$-values below the significance level $\alpha=0.05$ calculated based on $2000$ samples of $I=500$ independent pairs.
Time-specific Overall
Scenario $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $M_{\mathcal{T}}$ $PW$
1) No effect 0.052 0.047 0.052 0.055 0.050 0.049 0.056
2) PH 0.782 0.939 0.963 0.979 0.982 0.985 0.978
3) Early-div 0.865 0.943 0.937 0.873 0.758 0.954 0.957
4) Crossing 0.875 0.770 0.296 0.017 0.000 0.798 0.375
5) Late-div 0.124 0.345 0.657 0.880 0.970 0.927 0.610
4.2.   Design Sensitivities
We now use the explicit formulas (8) and (9) to asymptotically evaluate the powers of sensitivity analyses using $T_{\tau}$ and $M_{\mathcal{T}}$ via design sensitivities. Since the formulas are valid under random censoring, we consider the hazard model $h_{c}(\tau)=\lambda/b$ for $c_{ij}$, where $\lambda=0.2$ and $b>1$ is a constant for adjusting the censoring rate. Except this, the remaining simulation settings are the same as those described in subsection 4.1. For each the four alternative scenarios 2-5), $b$ is adjusted separately to achieve moderate (approximately $25\%$) non-administrative censoring. Under each alternative scenario, we calculate design sensitivities for $T_{\tau}$, $\tau=1,2,3,4,5$ and $M_{\mathcal{T}}$ with $\mathcal{T}=\{1,2,3,4,5\}$ by the following two steps: we first generate a single sample of $I=100,000$ independent paired sets of censored survival times and then use the Monte Carlo method to approximate the expectations involved in the formulas (8) and (9). The simulation results of design sensitivity are given in Table 2. Generally, higher design sensitivity implies higher power of sensitivity analysis (Zhang et al., 2024). Under the PH scenario, the design sensitivity of $T_{\tau}$ slightly increases as $\tau$ increases from $1$ to $5$. Under the three non-PH scenarios 3-5), the design sensitivities of the time-specific tests are relatively large at times with strong effects, indicating greater robustness to unmeasured confounding. Under the Crossing scenario, the adverse effect at $\tau=4$ and $5$ leads the design sensitivities of $T_{4}$ and $T_{5}$ to be less than $1$.
Table 2: Design sensitivities calculated based on a single sample of $I=100,000$ independent pairs under respective alternative scenario with random censoring.
Time-specific Overall
Scenario $\tilde{\Gamma}_{1}$ $\tilde{\Gamma}_{2}$ $\tilde{\Gamma}_{3}$ $\tilde{\Gamma}_{4}$ $\tilde{\Gamma}_{5}$ $\tilde{\Gamma}$
2) PH 1.491 1.530 1.549 1.560 1.568 1.567
3) Early-div 1.557 1.524 1.468 1.394 1.325 1.465
4) Crossing 1.574 1.371 1.156 $<$ 1 $<$ 1 1.398
5) Late-div 1.070 1.160 1.271 1.399 1.574 1.524
5.   Application
Following Kim et al. (2016), social engagement is measured using five baseline variables: gathering with friends, involvement in charity, leisure, culture or sports activity, religious meeting, and participation in alumni or hometown event. The variables are summed, with a higher total indicating more active social engagement. We focus on two groups: 1) socially active individuals with a total engagement score greater than 6 and 2) socially inactive individuals with a total engagement score of 3 or less. We form a treated group of 1474 socially active individuals and match them to a control group of 1474 socially inactive individuals. Matching variables include age, sex, education level, residential region, marital status, income, economic activity status, alcohol use, smoking history, depression, self-reported health status, and number of chronic diseases. The matching results are summarized in Table 3, which shows that, for the measured confounders, the absolute standardized mean differences (SMDs) are all less than 0.1.
Table 3: Confounder balance in 1474 matched treated-control pairs. We report the mean values of confounders and standardized mean differences (SMDs) before and after matching.
Before Matching After Matching
Confounder Treated Control SMD Treated Control SMD
Age 59.17 62.30 -0.305 59.17 59.18 -0.001
Age $\geq$ 65 0.315 0.426 -0.237 0.315 0.315 0.000
Male 0.423 0.442 -0.037 0.423 0.423 0.000
Education
      $\leq$ Elementary school 0.308 0.519 -0.457 0.308 0.310 -0.004
      Middle school 0.134 0.172 -0.111 0.134 0.124 0.030
      High school 0.341 0.241 0.211 0.341 0.386 -0.096
      $\geq$ College 0.215 0.068 0.357 0.215 0.179 0.088
      Missing 0.002 0.000 0.040 0.002 0.001 0.030
Urban Resident 0.832 0.794 0.101 0.832 0.827 0.013
Married 0.830 0.776 0.144 0.830 0.828 0.004
Income
      Low 0.200 0.265 -0.163 0.200 0.214 -0.034
      Middle 0.246 0.259 -0.029 0.246 0.242 0.011
      High 0.305 0.209 0.209 0.305 0.305 0.000
      Missing 0.249 0.267 -0.043 0.249 0.240 0.020
Economic Activity 0.406 0.359 0.094 0.406 0.409 -0.007
Alcohol use 0.428 0.447 -0.038 0.428 0.432 -0.008
Smoker 0.244 0.308 -0.149 0.244 0.254 -0.024
Depression 0.097 0.130 -0.112 0.097 0.087 0.034
Self-reported health status
      Bad 0.177 0.337 -0.418 0.177 0.189 -0.032
      Moderate 0.460 0.468 -0.017 0.460 0.467 -0.015
      Good 0.363 0.195 0.349 0.363 0.343 0.041
No. chronic diseases
      0 0.540 0.510 0.060 0.540 0.547 -0.015
      1 0.298 0.291 0.014 0.298 0.286 0.025
      $\geq$ 2 0.162 0.199 -0.099 0.162 0.166 -0.011
We apply our methods to a matched cohort study on how social engagement impacts later survival over time. We assess whether social engagement reported at the time of the baseline KLoSA survey is associated with a longer life, and identify the specific period during which this effect is significant. We look at the outcome of either survival or censoring time observed over the 16-year follow-up period between 2006-2022. Considering that the KLoSA is a biennial survey, we test $H_{0}(\tau)$ at $\tau=2,4,6,8,10,12,\ \text{and}\ 14$ follow-up years using $T_{\tau}$, and apply our overall testing procedure to test $H_{0}$ using $M_{\mathcal{T}}$ with $\mathcal{T}=\{2,4,6,8,10,12,14\}$. In addition to $M_{\mathcal{T}}$, we consider the Prentice-Wilcoxon statistic ($PW$) to test $H_{0}$.
Table 4 shows the results of the time-specific tests and overall test for various $\Gamma$ values. The upper bounds of $p$-values are reported. We find that, in the absence of unmeasured bias ($\Gamma=1$), the $p$-values for all time-specific tests fall below $\alpha=0.05$. This suggests that the impact of social engagement in 2006 on survival in the cohort sustained through 2022, assuming no unmeasured confounding. As $\Gamma$ increases, we can see that the worst-case $p$-values for the time-specific tests vary across given time points. The treatment effects at $\tau=2,4,6,8$ are quite strong even under larger values of $\Gamma$. For instance, we can reject $H_{0}(4)$ up to $\Gamma=1.6$ and reject $H_{0}(8)$ up to $\Gamma=1.3$. However, the effects at later time points can be comparatively weak and the $p$-values rapidly approach to 1 as $\Gamma$ increases. These effects can be explained if there is a small amount of unmeasured bias. This observed pattern is similar to the scenario of Early-div. in the simulation studies. We can see that the effect of social engagement on mortality is large in the early part of the study period. In this Early-div. scenario, there was a small difference between $M_{\tau}$ and PW with respect to power of test. However, in the sensitivity analysis, we can see that the conclusion obtained from $M_{\tau}$ is more robust to unmeasured bias. Regarding the overall tests, $M_{\mathcal{T}}$ leads to rejection of $H_{0}$ for $\Gamma\leq 1.2$ and $PW$ lead to rejection of $H_{0}$ for $\Gamma\leq 1.1$. We also found that our proposed method can be effectively integrated with PW. The results of this combined approach are presented in the column of the table. These findings are nearly identical to those obtained using $M_{\tau}$ alone, indicating that the addition of PW does not enhance performance. Surprisingly, there is a negligible contribution to multiple corrections by incorporating PW. In summary, we conclude that the treatment effect is large at the first 8 years but gets smaller as time goes on. Our proposed method can discover and confirm this pattern in a systematic approach, and the conclusion is much robust to unmeasured confounding.
Table 4: Sensitivity analysis for the matched KLoSA sample.
Time-specific Overall
Bias $T_{2}$ $T_{4}$ $T_{6}$ $T_{8}$ $T_{10}$ $T_{12}$ $T_{14}$ $M_{\mathcal{T}}$ $PW$ $M_{\mathcal{T}}+PW$
$\Gamma=1$ 0.001 0.000 0.000 0.000 0.005 0.013 0.037 0.000 0.003 0.000
$\Gamma=1.1$ 0.003 0.000 0.002 0.000 0.053 0.128 0.273 0.002 0.039 0.002
$\Gamma=1.15$ 0.005 0.000 0.004 0.002 0.120 0.262 0.480 0.009 0.091 0.009
$\Gamma=1.2$ 0.007 0.001 0.010 0.005 0.227 0.438 0.684 0.030 0.179 0.031
$\Gamma=1.25$ 0.011 0.001 0.020 0.012 0.366 0.620 0.838 0.082 0.302 0.084
$\Gamma=1.3$ 0.016 0.002 0.036 0.027 0.518 0.773 0.930 0.179 0.445 0.182
$\Gamma=1.35$ 0.021 0.004 0.061 0.054 0.663 0.880 0.974 0.321 0.591 0.324
$\Gamma=1.4$ 0.029 0.007 0.096 0.094 0.783 0.943 0.992 0.490 0.720 0.493
$\Gamma=1.45$ 0.038 0.012 0.142 0.152 0.871 0.976 0.998 0.656 0.823 0.658
$\Gamma=1.5$ 0.049 0.019 0.199 0.226 0.929 0.991 0.999 0.792 0.895 0.794
$\Gamma=1.55$ 0.061 0.028 0.264 0.314 0.964 0.997 1.000 0.888 0.943 0.888
$\Gamma=1.6$ 0.075 0.041 0.337 0.411 0.983 0.999 1.000 0.945 0.970 0.946
$\Gamma=1.65$ 0.092 0.057 0.414 0.510 0.992 1.000 1.000 0.976 0.986 0.976
We are also interested in testing the time-specific hypothesis $H_{0}(\tau)$. As discussed in Section 3.4, we apply the closed testing procedure to examine each of $H_{0}(\tau)$. Table 5 shows the maximum value of the $p$-value upper bound, $\max_{\mathcal{I}_{A}}p(\mathcal{I}_{A})$. For instance, when $A=\{4\}$, the results are the same as the results of $M_{\tau}$ in Table 4. This is because the maximum statistic $M_{\tau}$ is obtained from the time-specific test statistic $T_{4}$. $H_{0}(4)$ was rejected up to $\Gamma=1.6$, but after the multiple testing correction, it is now rejected up to $\Gamma=1.2$.
Table 5: Maximum of the $p$-value upper bounds, $\max_{\mathcal{I}_{A}}p(\mathcal{I}_{A})$, in the closed testing procedure.
Time-specific
Bias $H_{0}(2)$ $H_{0}(4)$ $H_{0}(6)$ $H_{0}(8)$ $H_{0}(10)$ $H_{0}(12)$ $H_{0}(14)$
$\Gamma=1$ 0.003 0.000 0.001 0.000 0.011 0.021 0.037
$\Gamma=1.1$ 0.045 0.002 0.016 0.003 0.121 0.187 0.273
$\Gamma=1.15$ 0.120 0.009 0.051 0.014 0.262 0.365 0.480
$\Gamma=1.2$ 0.252 0.030 0.129 0.046 0.451 0.570 0.684
$\Gamma=1.25$ 0.430 0.082 0.260 0.116 0.645 0.751 0.838
6.   Discussion
In this work, we developed methods for evaluating a treatment effect on a time-to-event outcome in matched observational studies. Our contribution is manifold. First, we proposed a time-specific test that assesses a treatment effect within a specific duration. Sequential use of this test allows for a granular analysis of a treatment impact over time, which traditional approaches cannot achieve. This strategy will be particularly useful in studies aiming to demonstrate a treatment’s efficacy within a set time or to understand how a treatment impacts a time-to-event outcome over time. Second. We introduced an overall testing procedure that jointly uses time-specific tests from distinct analysis times and demonstrated that this approach can accommodate far more complex treatment effects than PH. From our results, we anticipate a relatively enhanced applicability of our overall test in varying alternatives compared to traditional methods. Third, we extended our methods to general matching designs by developing corresponding sensitivity analysis methods for observational studies.
We suggest three future research directions based on our findings. First, our method can be improved by finding more effective time-specific scores than pseudo-observations. Particularly at early time points, where events rarely occur for most individuals, many scores are uninformative. If there is an alternative approach of computing scores during these early periods, any initial difference can be more easily detected. Second, the inequality (6) may not be sufficiently precise when $\Gamma>1$. A more computationally intensive strategy could be explored to improve its bound, as suggested by Fogarty and Small (2016). Implementing such a method requires effectively integrating the correlation structure of the test statistics. Lastly, in this study, the time points were predetermined based on biennial data collection. A more interesting approach would involve developing a data-driven approach that adaptively selects the number and location of time points to analyze the effect. This could lead to more tailored and potentially insightful findings.
Data Availability
The data used in this study are from the Korean Longitudinal Study of Aging (KLoSA) and are publicly available through the Korea Employment Information Service (KEIS) website (https://survey.keis.or.kr/index.jsp)
Appendix: Technical Proofs
Proof of Theorem 1
Theorem 1 follows from the multivariate central limit theorem by Fabian and Hannan (1985, p. 164), taking their $n$ to be $I$ and their $\bm{X}_{nj}$ to be $(d_{j}(\tau_{1})/\sigma_{\tau_{1}},\dots,d_{j}(\tau_{L})/\sigma_{\tau_{L}})^{T}V_{j}$. Define $\bm{e}_{l}$ to be the unit vector in $\mathbb{R}^{L}$ with its $l$th entry being $1$. By the argument in the central limit theorem, it is enough to verify that, for each $\bm{e}_{l}$, the triangular array $\bm{e}_{l}^{T}\bm{X}_{n1},\dots,\bm{e}_{l}^{T}\bm{X}_{nn}$, $n\geq 1$ satisfy the Lindberg condition,
$\begin{split}\lim_{n\to\infty}\sum_{j=1}^{n}E\left[(\bm{e}_{l}^{T}\bm{X}_{nj})^{2}I(|\bm{e}_{l}^{T}\bm{X}_{nj}|>\epsilon)|\mathcal{F},\mathcal{V}\right]=0\ \text{for any}\ \epsilon>0.\end{split}$
For each $\tau_{l}\in\mathcal{T}$ and a given $\epsilon>0$,
$\begin{split}&\sum_{j=1}^{n}E\left[(\bm{e}_{l}^{T}\bm{X}_{nj})^{2}I(|\bm{e}_{l}^{T}\bm{X}_{nj}|>\epsilon)|\mathcal{F},\mathcal{V}\right]\\ &=\sum_{j=1}^{n}E\left[\left(d_{j}(\tau_{l})V_{j}/\sigma_{\tau_{l}}\right)^{2}I\left(\left|d_{j}(\tau_{l})V_{j}/\sigma_{\tau_{l}}\right|>\epsilon\right)\bigg{|}\mathcal{F},\mathcal{V}\right]\\[5.0pt] &=\sum_{j=1}^{n}(d_{j}(\tau_{l})/\sigma_{\tau_{l}})^{2}E\left[V_{j}^{2}I\left(d_{j}^{2}(\tau_{l})V_{j}^{2}/\sigma_{\tau_{l}}^{2}>\epsilon^{2}\right)\bigg{|}\mathcal{F},\mathcal{V}\right]\\[5.0pt] &\leq E\left[V_{1}^{2}I\left(\max_{j\in\{1,\dots,n\}}\frac{d_{j}^{2}(\tau_{l})}{\sum_{j=1}^{n}d_{j}^{2}(\tau_{l})}V_{1}^{2}>\epsilon^{2}\right)\bigg{|}\mathcal{F},\mathcal{V}\right]\\[10.00002pt] &\leq E\left[V_{1}^{2}I\left(\max_{\tau_{l}\in\mathcal{T}}\left\{\max_{j\in\{1,\dots,n\}}\frac{d_{i}^{2}(\tau_{l})}{\sum_{j=1}^{n}d_{j}^{2}(\tau_{l})}\right\}V_{1}^{2}>\epsilon^{2}\right)\bigg{|}\mathcal{F},\mathcal{V}\right],\end{split}$
where the third relation holds under randomized treatment assignments. Using the condition in Theorem 1, the last line converges to 00 as $n\rightarrow\infty$ by the dominated convergence theorem, proving Theorem 1. Strictly speaking, as $n\rightarrow\infty$ in this proof, the notation should refer to sequences of $\mathcal{F}_{n}$ and $\mathcal{V}_{n}$, changing with $n$, but to avoid cumbersome notation, this was not indicated explicitly. $\square$
Proof of Theorem 2
The proof of Theorem 2 is parallel to that of Theorem 1, taking $n$ to be $I$ and $\bm{X}_{nj}$ to be $((d_{j}(\tau_{1})V_{\tau_{1},j}^{+}-\{(\Gamma-1)/(1+\Gamma)\}|d_{j}(\tau_{1})|)/\sigma_{\tau_{1}}^{+},\dots,(d_{j}(\tau_{L})V_{\tau_{L},j}^{+}-\{(\Gamma-1)/(1+\Gamma)\}|d_{j}(\tau_{L})|)/\sigma_{\tau_{L}}^{+})^{T}$ where $V_{\tau_{l},j}^{+}$ denotes the distribution of $V_{j}$ when $w_{j}=w_{\tau_{l},j}^{+}$. $\square$
Proof of Theorem 3
By the arguments in Section 2.4, under the conditions in Lemma 1, we can rewrite $d_{i}(\tau)V_{i}$ as $-(\dot{\psi}(Y_{i1},\delta_{i1},\tau)-\dot{\psi}(Y_{i2},\delta_{i2},\tau))V_{i}+o_{P}(1)$. As $I\to\infty$, using that $\dot{\psi}(\cdot;\tau)$ is bounded, the numerator of $D_{\tau,I}$ converges in probability to $\{E(d_{i}(\tau)V_{i})-[(\Gamma-1)/(1+\Gamma)]E(|d_{i}(\tau)|)\}/\sqrt{E(d_{i}^{2}(\tau))}$ while the denominator of $D_{\tau,I}$ tends to 00. The formula (8) is then obtained by rearranging the equation $\{E(d_{i}(\tau)V_{i})-[(\Gamma-1)/(1+\Gamma)]E(|d_{i}(\tau)|)\}/\sqrt{E(d_{i}^{2}(\tau))}=0$. $\square$
Proof of Theorem 4
Consider the limiting behavior of the minimum among
$\frac{{\displaystyle\max_{\tau_{l}\in\mathcal{T}}}\frac{\sum_{i=1}^{I}d_{i}(\tau_{l})V_{i}/I}{\sqrt{\sum_{i=1}^{I}d_{i}^{2}(\tau_{l})/I}}-\frac{\Gamma-1}{1+\Gamma}\frac{\sum_{i=1}^{I}|d_{i}(\tau_{l})|/I}{\sqrt{\sum_{i=1}^{I}d_{i}^{2}(\tau_{l})/I}}}{\sqrt{\frac{4\Gamma}{(1+\Gamma)^{2}}/I}}=D_{\tau_{l},I}^{*},\tau_{l}\in\mathcal{T},$
as $I\to\infty$. By a similar argument to the proof of Theorem 3, as $I\to\infty$, the numerator of $\min_{\tau_{l}\in\mathcal{T}}D_{\tau_{l},I}^{*}$ converges in probability to $\max_{\tau_{l}\in\mathcal{T}}E(d_{i}(\tau_{l})V_{i})/\sqrt{E(d_{i}^{2}(\tau_{l}))}-[(\Gamma-1)/(1+\Gamma)]\max_{\tau_{l}\in\mathcal{T}}E(|d_{i}(\tau_{l})|)/\sqrt{E(d_{i}^{2}(\tau_{l}))}$ while the denominator of $\min_{\tau_{l}\in\mathcal{T}}D_{\tau_{l},I}^{*}$ tends to 00. Observing that
$\Pr\left(\chi_{l}\leq\min_{\tau_{l}\in\mathcal{T}}D_{\tau_{l},I}^{*},l=1,\dots,L\right)\leq\Lambda_{\mathcal{T},I}\leq\Pr\left(\chi_{1}\leq\min_{\tau_{l}\in\mathcal{T}}D_{\tau_{l},I}^{*}\right),$
$\Lambda_{\mathcal{T},I}$ tends to $1$ for $\Gamma<\tilde{\Gamma}_{\mathcal{T}}$ and to 00 for $\Gamma>\tilde{\Gamma}_{\mathcal{T}}$ as $I\to\infty$. $\square$